# Parametrically enhancing sensor sensitivity at an exceptional point

### 08 Dec 2023 v1

@@ I Introduction
Exceptional Points (EPs), Non-Hermitian degeneracy, are known in physical systems for their interesting counter-intuitive features and intriguing effects. Among the fascinating properties of EPs, we can mention nonreciprocity [1, 2], topological transport phenomena [3, 4], stopping light [5], loss-induced suppression and revival of lasing, pump-induced lasing death and unidirectional invisibility (see [6] and references therein). Recently, EPs have been used to enhance sensor sensitivity and versatile physical systems were proposed for this purpose including optical/photonic [7, 8], microwave [9], plasmonics [10, 11], and mechanic/acoustic [12, 13]. Coupled optomechanical systems, which result from a coupling between an electromagnetic field and a mechanical object [14, 15], have also been used to engineer EPs owing to their ability to induce gain and losses depending on their sideband pumping mechanism. Exceptional points engineering in optomechanics have led to a number of interesting phenomena such as mass sensing improvement [16, 17], topological energy transfer [18, 19], and nonlinearly induced collective phenomena [20, 21] among others.
@@
Exceptional point sensing is an approach which provides better sensitivity compared to conventional sensing schemes where the perturbation induces a shift in the frequency. The resulting shift/signal is generally weak and needs a pre-amplification process in order to be detected and analyzed [22]. For the Non-Hermitian sensors, however, the signal is self-amplified owing to the abrupt singularity and the topological feature at the EP, leading to a giant sensitivity enhancement without a pre-amplification requirement of the detected signal. This new sensing scheme has recently attracted attention as a means to enhance the responsiveness of sensors, and has started to be practically implemented. For instance, exceptional point sensing has been used to monitor internal physiological states via wireless interrogation techniques [23], and for an efficient wireless power transfer [24, 25]. Beside these interesting sensing applications probed at EPs, the implementation of EP-sensors is still in its infancy stage and it seems to rise up scepticism regarding noise amplification, which induces a degradation of the sensor’s performance [26]. This noise issue can be of a fundamental nature owing to the eigenbasis collapse or of a technical nature related to the amplification mechanisms used for the realization of EPs. This question has been a bottleneck of EP-sensors, however, appropriate techniques have started to be used to handle and suppress this noise issue [27, 28, 29]. Therefore, tremendous research activities are still going on to tackle and to get rid of this limitation [30, 31, 32, 33], and this includes nonlinear phenomena [34, 35].
@@
The current work introduces a new mechanism to enhance the sensitivity of EP-sensors, which is based on a mechanical parametric driving effect. Our proposal consists of two coupled optomechanical systems, where the mechanical resonators are mechanically coupled. The parametric driving comes from a nonlinear modulation of the spring constants of the mechanical resonators [36, 37]. We investigated two sensing schemes. The first one is based on the splitting [16] of the EP and the second is related to the shifting [38, 39] of the EP. When parametric driving is switched on, we found that i) the sensitivity is greatly enhanced, ii) the EP is shifted towards low frequencies, and iii) the strength of the optical driving field corresponding to the EP increases. Moreover, it has been observed that the overall sensitivity enhancement in the shifting scenario is weak compared to the enhancement captured with the splitting scheme. Furthermore, the output spectra and transmissions of the optical cavities have been investigated under some system parameters. The effect of perturbations, the parametric driving impact and thermal noise on the EP were analyzed via the spectra. While the effect of the other parameters is confirmed as aforementioned, it has been found that thermal noise amplifies the output signal, leading to better sensitivity. This work shows how sensing based on the splitting of the EP outperforms sensors operating in a scenario where the EP is shifted. Furthermore, it also paves a way towards improvement of sensing through nonlinear effects in optomechanics [40, 41, 42].
@@ II Model and dynamical equations
Figure 1: Sketch of our benchmark system. Cavity $1$ ($2$) is driven with red (blue) electromagnetic field to generate losses (gain). The sprint constants of the involved mechanical resonators are modulated to induce parametric driving [36, 37].
@@
In Eq. 1, $H_{\rm{0M}}$, $H_{\rm{PA}}$, $H_{\rm{int}}$, and $H_{\rm{drive}}$ are the Hamiltonians describing the optomechanical cavities, the mechanical parametric amplification (MPA), the mechanical interactions and the driving fields, respectively. The annihilation (creation) operators related to the mechanical resonators and optical fields are $b_{j}$ ($b^{\dagger}_{j}$) and $a_{j}$ ($a_{j}^{\dagger}$), respectively. The other parameters are the mechanical ($\omega_{j}$) and cavity ($\omega_{c}^{j}$) frequencies, the optomechanical coupling ($g_{j}$), the phonon hopping coupling rate ($J_{m}$) and the driving amplitude ($\rm{E_{j}}$). The parametric driving amplitude $\chi_{j}$ comes from the modulation of the spring constants of the mechanical resonators at frequency $2\omega_{d}$ and phase $\phi_{d}$ [36, 37]. For the sake of simplicity and without loss of generality, we assume that the optical cavities are driven by the same amount of driving strength and are degenerated, so that from now on we consider $\rm{E_{1}}=\rm{E_{2}}=\rm{E}$, $\omega^{1}_{c}=\omega^{2}_{c}=\omega_{c}$, and $g_{1}=g_{2}=g$. Throughout the work, we also assume $\phi_{d}=0$ and $\chi_{1}=\chi_{2}=\chi$. In the rotating wave approximation (RWA) and in the frame rotating at $\omega_{p}^{j}+\omega_{d}$, where $\omega_{p}^{j}$ is the $j^{th}$ electromagnetic driving frequency, the above Hamiltonian yields:
@@ III Sensitivity through a splitting and shifting of an exceptional point
where $\sigma=\sqrt{16\tilde{J}_{m}^{2}+(2\delta{\Delta_{m}}-i\delta\rm{\Gamma_{eff}})^{2}}$ with $\delta{\Delta_{m}}=\Delta_{m}^{1}-\Delta_{m}^{2}$ and $\delta\rm{\Gamma_{eff}}=\rm{\Gamma_{eff}}^{1}-\rm{\Gamma_{eff}}^{2}$. The eigenvalues of our system are given by Eq. 15, and are displayed in Fig. 2, where the eigenfrequencies ($\omega_{\pm}=\operatorname{Re}(\lambda_{\pm})$) and the dissipations ($\gamma_{\pm}=\operatorname{Im}(\lambda_{\pm})$) are shown in Fig. 2(a,c) and Fig. 2(b,d), respectively. In this figure, the parametric driving is off ($\chi=0$) and the solid (dashed) line depicts the case before (after) the perturbation of the system. In Fig. 2(a,b) the perturbation induces a splitting of the $\rm{EP}$ while in Fig. 2(c,d), it instead induces a shift of the $\rm{EP}$. In the former case, any mass deposition ($\delta m$) acts as a perturbation on the system, and induces a frequency shift ($\delta\omega$) through the relationship, $\delta m=\mathcal{R}^{-1}\delta\omega$ where $\mathcal{R}=\frac{\omega_{m}}{2m}$ is the mass responsibility and $\rm{m}$ is the mass of the resonator supporting the deposition. Therefore, this mass deposition leads to a frequency shift that lifts up the $\rm{EP}$-degeneracy, resulting in a sensitivity enhancement at the $\rm{EP}$ (see double arrow in Fig. 2(a,b)). In the latter case, the perturbation induces a shift in the dissipation instead, which can be seen as a dissipation mismatch ($\delta\gamma=|\gamma_{1}-\gamma_{2}|$) between the mechanical resonators [38, 39]. This perturbation can also be thought of as induced by a mismatch between optical damping ($\delta\gamma=|\Gamma_{1}-\Gamma_{2}|$), and may originate from optical fluctuations within the cavities. This scenario also leads to a sensing scheme as shown by the double arrow in Fig. 2(c,d). A general observation from Fig. 2 is that after a splitting due to a mass deposition (Fig. 2(a,b)), the $\rm{EP}$ is destroyed and that limits the ability of the system to perform multiple sensing scenarios in a time sequence. Indeed, the enhanced performance is deteriorated after each sensing process since every next detection is conducted on the basis of being destroyed by the previous detection. In order to preserve $\rm{EP}$ in the parameter space after a perturbation, it is preferable to shift the $\rm{EP}$ instead of lifting it up. Such a sensor based on shifting the $\rm{EP}$ (see Fig. 2(c,d)) will therefore meet the condition of a non-demolition sensing mechanism. With a fine adjustment of gain and loss in the system, we expect a linewidth preservation of the eigenstates after optical perturbation in the sensing scheme based on shifting the $\rm{EP}$. This may improve the performance of such sensing scenarios in the perspective of experiments related to the precision of dispersion measurement [39]. However, we stress that the sensitivity of this novel sensing scheme is somehow weak compared to the splitting scheme (compare the double arrows in Fig. 2 (a) and (c)).
@@
with $\nu=\frac{1}{2}\left(\frac{\chi_{2}^{2}}{\omega_{m}}-\frac{\chi_{1}^{2}}{\omega_{m}}\right)-\frac{\omega_{m}\delta\omega}{\Delta_{m}^{2}}$ and $\mu=-i\delta\Gamma_{\rm{eff}}$. In these equations (Eq. 17 and Eq. 18), all scenarios of perturbations allowed in our system have been considered. By considering that there is no parametric driving ($\chi_{j}=0$), and that there is no perturbation of the dissipation ($\delta\gamma=0$), the system reduce to the one investigated in [16], and the expressions in Eq. 17 and Eq. 18 reduce to $\Delta\omega^{\epsilon}=\frac{\delta\omega_{2}}{2}$ and $\sigma_{\rm{EP}}^{\epsilon}=(1+i)\sqrt{\frac{\delta\Gamma_{\rm{eff}}\delta\omega_{2}}{8}}$ as expected. This sensing scenario is based on a splitting at the $\rm{EP}$ as it can be seen in Fig. 2(a,b). In that case, the amplification factor is given by:
@@ Acknowledgments
This work has been carried out under the Iso-Lomso Fellowship at Stellenbosch Institute for Advanced Study (STIAS), Wallenberg Research Centre at Stellenbosch University, Stellenbosch 7600, South Africa.
P. Djorwe acknowledges the receipt of a grant from the APS-EPS-FECS-ICTP Travel Award Fellowship Programme (ATAP), Trieste, Italy.
@@ Appendix A Effective mechanical Hamiltonian
$\small{\begin{cases}\delta\tilde{\dot{\alpha_{j}}}&=-\frac{\kappa}{2}\delta\tilde{\alpha_{j}}+i\tilde{G_{j}}(\delta\tilde{\beta_{j}}+\delta\tilde{\beta_{j}}^{\dagger}e^{2i\Delta_{m}^{j}t})+\sqrt{\kappa}\delta\tilde{\alpha^{in}_{j}},\\ \delta\tilde{\dot{\beta_{j}}}&=\rm{-\frac{\gamma_{m}}{2}\delta\tilde{\beta_{j}}+i\tilde{J}_{m}\delta\tilde{\beta}_{3-j}+i\tilde{G_{j}}(\delta\tilde{\alpha_{j}}^{\dagger}e^{2i\tilde{\Delta}_{j}t}+\delta\tilde{\alpha_{j}})+\sqrt{\gamma_{m}}\delta\tilde{\beta}^{in}_{j}},\end{cases}}$ (35)

### 12 Dec 2023 v2

@@ I Introduction
Exceptional Points (EPs), Non-Hermitian degeneracy, are known in physical systems for their interesting counter-intuitive features and intriguing effects. Among the fascinating properties of EPs, we can mention nonreciprocity Thomas et al. (2016); Lau and Clerk (2018), topological transport phenomena Xu et al. (2023a, b), stopping light Goldzak et al. (2018), loss-induced suppression and revival of lasing, pump-induced lasing death and unidirectional invisibility (see Peng et al. (2014) and references therein). Recently, EPs have been used to enhance sensor sensitivity and versatile physical systems were proposed for this purpose including optical/photonic Hodaei et al. (2017); Chen et al. (2017), microwave Nada et al. (2018), plasmonics Park et al. (2020); Jiang et al. (2022), and mechanic/acoustic Rosa et al. (2021); Cai et al. (2022). Coupled optomechanical systems, which result from a coupling between an electromagnetic field and a mechanical object Djorwé et al. (2013a); Aspelmeyer et al. (2014), have also been used to engineer EPs owing to their ability to induce gain and losses depending on their sideband pumping mechanism. Exceptional points engineering in optomechanics have led to a number of interesting phenomena such as mass sensing improvement Djorwe et al. (2019a); Tchounda et al. (2023), topological energy transfer Chen et al. (2023); Xu et al. (2016), and nonlinearly induced collective phenomena Djorwe et al. (2018); Djorwé et al. (2020) among others.
@@
Exceptional point sensing is an approach which provides better sensitivity compared to conventional sensing schemes where the perturbation induces a shift in the frequency. The resulting shift/signal is generally weak and needs a pre-amplification process in order to be detected and analyzed Zhu et al. (2021). For the Non-Hermitian sensors, however, the signal is self-amplified owing to the abrupt singularity and the topological feature at the EP, leading to a giant sensitivity enhancement without a pre-amplification requirement of the detected signal. This new sensing scheme has recently attracted attention as a means to enhance the responsiveness of sensors, and has started to be practically implemented. For instance, exceptional point sensing has been used to monitor internal physiological states via wireless interrogation techniques Chen and El-Ganainy (2019), and for an efficient wireless power transfer Assawaworrarit et al. (2017); Song et al. (2021). Beside these interesting sensing applications probed at EPs, the implementation of EP-sensors is still in its infancy stage and it seems to rise up scepticism regarding noise amplification, which induces a degradation of the sensor’s performance Anderson et al. (2023). This noise issue can be of a fundamental nature owing to the eigenbasis collapse or of a technical nature related to the amplification mechanisms used for the realization of EPs. This question has been a bottleneck of EP-sensors, however, appropriate techniques have started to be used to handle and suppress this noise issue Kononchuk et al. (2022); Suntharalingam et al. (2023); Tuxbury et al. (2022). Therefore, tremendous research activities are still going on to tackle and to get rid of this limitation Chen and Zhao (2020); Liu et al. (2021); Wu et al. (2023); Khanbekyan (2023), and this includes nonlinear phenomena Li et al. (2023); Bai et al. (2022).
@@
The current work introduces a new mechanism to enhance the sensitivity of EP-sensors, which is based on a mechanical parametric driving effect. Our proposal consists of two coupled optomechanical systems, where the mechanical resonators are mechanically coupled. The parametric driving comes from a nonlinear modulation of the spring constants of the mechanical resonators Lemonde et al. (2016); Bothner et al. (2020). We investigated two sensing schemes. The first one is based on the splitting Djorwe et al. (2019a) of the EP and the second is related to the shifting Wang et al. (2019); Mao et al. (2023) of the EP. When parametric driving is switched on, we found that i) the sensitivity is greatly enhanced, ii) the EP is shifted towards low frequencies, and iii) the strength of the optical driving field corresponding to the EP increases. Moreover, it has been observed that the overall sensitivity enhancement in the shifting scenario is weak compared to the enhancement captured with the splitting scheme. Furthermore, the output spectra and transmissions of the optical cavities have been investigated under some system parameters. The effect of perturbations, the parametric driving impact and thermal noise on the EP were analyzed via the spectra. While the effect of the other parameters is confirmed as aforementioned, it has been found that thermal noise amplifies the output signal, leading to better sensitivity. This work shows how sensing based on the splitting of the EP outperforms sensors operating in a scenario where the EP is shifted. Furthermore, it also paves a way towards improvement of sensing through nonlinear effects in optomechanics Djorwé et al. (2013b); Leijssen et al. (2017); Djorwe et al. (2019b).
@@ II Model and dynamical equations
Figure 1: Sketch of our benchmark system. Cavity $1$ ($2$) is driven with red (blue) electromagnetic field to generate losses (gain). The sprint constants of the involved mechanical resonators are modulated to induce parametric driving Lemonde et al. (2016); Bothner et al. (2020).
@@
In Eq. 1, $H_{\rm{0M}}$, $H_{\rm{PA}}$, $H_{\rm{int}}$, and $H_{\rm{drive}}$ are the Hamiltonians describing the optomechanical cavities, the mechanical parametric amplification (MPA), the mechanical interactions and the driving fields, respectively. The annihilation (creation) operators related to the mechanical resonators and optical fields are $b_{j}$ ($b^{\dagger}_{j}$) and $a_{j}$ ($a_{j}^{\dagger}$), respectively. The other parameters are the mechanical ($\omega_{j}$) and cavity ($\omega_{c}^{j}$) frequencies, the optomechanical coupling ($g_{j}$), the phonon hopping coupling rate ($J_{m}$) and the driving amplitude ($\rm{E_{j}}$). The parametric driving amplitude $\chi_{j}$ comes from the modulation of the spring constants of the mechanical resonators at frequency $2\omega_{d}$ and phase $\phi_{d}$ Lemonde et al. (2016); Bothner et al. (2020). For the sake of simplicity and without loss of generality, we assume that the optical cavities are driven by the same amount of driving strength and are degenerated, so that from now on we consider $\rm{E_{1}}=\rm{E_{2}}=\rm{E}$, $\omega^{1}_{c}=\omega^{2}_{c}=\omega_{c}$, and $g_{1}=g_{2}=g$. Throughout the work, we also assume $\phi_{d}=0$ and $\chi_{1}=\chi_{2}=\chi$. In the rotating wave approximation (RWA) and in the frame rotating at $\omega_{p}^{j}+\omega_{d}$, where $\omega_{p}^{j}$ is the $j^{th}$ electromagnetic driving frequency, the above Hamiltonian yields:
@@ III Sensitivity through a splitting and shifting of an exceptional point
where $\sigma=\sqrt{16\tilde{J}_{m}^{2}+(2\delta{\Delta_{m}}-i\delta\rm{\Gamma_{eff}})^{2}}$ with $\delta{\Delta_{m}}=\Delta_{m}^{1}-\Delta_{m}^{2}$ and $\delta\rm{\Gamma_{eff}}=\rm{\Gamma_{eff}}^{1}-\rm{\Gamma_{eff}}^{2}$. The eigenvalues of our system are given by Eq. 15, and are displayed in Fig. 2, where the eigenfrequencies ($\omega_{\pm}=\operatorname{Re}(\lambda_{\pm})$) and the dissipations ($\gamma_{\pm}=\operatorname{Im}(\lambda_{\pm})$) are shown in Fig. 2(a,c) and Fig. 2(b,d), respectively. In this figure, the parametric driving is off ($\chi=0$) and the solid (dashed) line depicts the case before (after) the perturbation of the system. In Fig. 2(a,b) the perturbation induces a splitting of the $\rm{EP}$ while in Fig. 2(c,d), it instead induces a shift of the $\rm{EP}$. In the former case, any mass deposition ($\delta m$) acts as a perturbation on the system, and induces a frequency shift ($\delta\omega$) through the relationship, $\delta m=\mathcal{R}^{-1}\delta\omega$ where $\mathcal{R}=\frac{\omega_{m}}{2m}$ is the mass responsibility and $\rm{m}$ is the mass of the resonator supporting the deposition. Therefore, this mass deposition leads to a frequency shift that lifts up the $\rm{EP}$-degeneracy, resulting in a sensitivity enhancement at the $\rm{EP}$ (see double arrow in Fig. 2(a,b)). In the latter case, the perturbation induces a shift in the dissipation instead, which can be seen as a dissipation mismatch ($\delta\gamma=|\gamma_{1}-\gamma_{2}|$) between the mechanical resonators Wang et al. (2019); Mao et al. (2023). This perturbation can also be thought of as induced by a mismatch between optical damping ($\delta\gamma=|\Gamma_{1}-\Gamma_{2}|$), and may originate from optical fluctuations within the cavities. This scenario also leads to a sensing scheme as shown by the double arrow in Fig. 2(c,d). A general observation from Fig. 2 is that after a splitting due to a mass deposition (Fig. 2(a,b)), the $\rm{EP}$ is destroyed and that limits the ability of the system to perform multiple sensing scenarios in a time sequence. Indeed, the enhanced performance is deteriorated after each sensing process since every next detection is conducted on the basis of being destroyed by the previous detection. In order to preserve $\rm{EP}$ in the parameter space after a perturbation, it is preferable to shift the $\rm{EP}$ instead of lifting it up. Such a sensor based on shifting the $\rm{EP}$ (see Fig. 2(c,d)) will therefore meet the condition of a non-demolition sensing mechanism. With a fine adjustment of gain and loss in the system, we expect a linewidth preservation of the eigenstates after optical perturbation in the sensing scheme based on shifting the $\rm{EP}$. This may improve the performance of such sensing scenarios in the perspective of experiments related to the precision of dispersion measurement Mao et al. (2023). However, we stress that the sensitivity of this novel sensing scheme is somehow weak compared to the splitting scheme (compare the double arrows in Fig. 2 (a) and (c)).
@@
with $\nu=\frac{1}{2}\left(\frac{\chi_{2}^{2}}{\omega_{m}}-\frac{\chi_{1}^{2}}{\omega_{m}}\right)-\frac{\omega_{m}\delta\omega}{\Delta_{m}^{2}}$ and $\mu=-i\delta\Gamma_{\rm{eff}}$. In these equations (Eq. 17 and Eq. 18), all scenarios of perturbations allowed in our system have been considered. By considering that there is no parametric driving ($\chi_{j}=0$), and that there is no perturbation of the dissipation ($\delta\gamma=0$), the system reduce to the one investigated in Djorwe et al. (2019a), and the expressions in Eq. 17 and Eq. 18 reduce to $\Delta\omega^{\epsilon}=\frac{\delta\omega_{2}}{2}$ and $\sigma_{\rm{EP}}^{\epsilon}=(1+i)\sqrt{\frac{\delta\Gamma_{\rm{eff}}\delta\omega_{2}}{8}}$ as expected. This sensing scenario is based on a splitting at the $\rm{EP}$ as it can be seen in Fig. 2(a,b). In that case, the amplification factor is given by:
@@ Acknowledgments
This work has been carried out under the Iso-Lomso Fellowship at Stellenbosch Institute for Advanced Study (STIAS), Wallenberg Research Centre at Stellenbosch University, Stellenbosch 7600, South Africa. P. Djorwe acknowledges the receipt of a grant from the APS-EPS-FECS-ICTP Travel Award Fellowship Programme (ATAP), Trieste, Italy. M. Asjad and D. Dutykh have been supported by the Khalifa University of Science and Technology under Award No. FSU- 2023-014.
@@ Appendix A Effective mechanical Hamiltonian
$\delta\tilde{\dot{\alpha_{j}}}$ $=-\frac{\kappa}{2}\delta\tilde{\alpha_{j}}+i\tilde{G_{j}}(\delta\tilde{\beta_{j}}+\delta\tilde{\beta_{j}}^{\dagger}e^{2i\Delta_{m}^{j}t})+\sqrt{\kappa}\delta\tilde{\alpha^{in}_{j}},$ (35a)
$\delta\tilde{\dot{\beta_{j}}}$ $=\rm{-\frac{\gamma_{m}}{2}\delta\tilde{\beta_{j}}+i\tilde{J}_{m}\delta\tilde{\beta}_{3-j}+i\tilde{G_{j}}(\delta\tilde{\alpha_{j}}^{\dagger}e^{2i\tilde{\Delta}_{j}t}+\delta\tilde{\alpha_{j}})+\sqrt{\gamma_{m}}\delta\tilde{\beta}^{in}_{j}},$ (35b)